Find the measure of an angle whose supplement measures 39 degrees more than twice its complement.
Answer by malakumar_kos@yahoo.com(315)


Find the measure of an angle whose supplement measures 39 degrees more than twice its complement.
Let the required angle be = x degrees
its supplement will be = 180-x
its copliment will be = 90-x
given supplement = 2(compliment)+39
180-x = 2(90-x)+39
2x-x = 180-180+39
x = 39 degrees